OpenStudy (anonymous):

solve the equation: the square root of x^2-3x+25=x+2

OpenStudy (anonymous):

set it =0 and use the quadratic formula.

OpenStudy (anonymous):

\left\{\left\{x\to 2-i \sqrt{19}\right\},\left\{x\to 2+i \sqrt{19}\right\}\right\}

OpenStudy (anonymous):

\[\left\{\left\{x\to 2-i \sqrt{19}\right\},\left\{x\to 2+i \sqrt{19}\right\}\right\}\]

OpenStudy (anonymous):

\[\sqrt{x ^{2}-3x+25}=x+2\]

OpenStudy (anonymous):

wow what are you doing? That doesn't look right at all.

OpenStudy (anonymous):

thats what the problem is , that s what i have to solve

OpenStudy (anonymous):

Do you know what the quadratic formula is?

OpenStudy (anonymous):

yeah but you dont use the quadratic formula for this problem

OpenStudy (anonymous):

you got to square both sides but after that i have no idea what to do

OpenStudy (anonymous):

Why can't you use the quadratic formula?

OpenStudy (anonymous):

because this is solving radical equations

OpenStudy (anonymous):

But this time you have another value of x aside from x^2

OpenStudy (anonymous):

\[(\sqrt{x ^{2}-3x+25})^{2}=(x+2)^{2}\]

OpenStudy (anonymous):

that problem is that you took a square root out of nowhere.

OpenStudy (anonymous):

This is the first step in solving it , but then i dont know what to do, this is how i have it in my notes and how it has it in the book

OpenStudy (anonymous):

a square cancels the square root on the left side

OpenStudy (anonymous):

But if you take the square root of one side you have to take the square root of the other side. So that squared thing is pointless.

OpenStudy (anonymous):

idk that s how the professor told us to do it

OpenStudy (anonymous):

Ok do that square root thing but do it to BOTH SIDES